What is the mass of silver bromide precipitated frromm 2.96g of iron bromide
1
Expert's answer
2019-11-21T06:25:49-0500

The reaction can be shown as follows:

FeBr3 + 3Ag+ = 3AgBr + Fe3+

From here, the mass of precipitated silver bromide equals:

m(AgBr) = m(FeBr3) × 3 × Mr(AgBr) / Mr (FeBr3) = 2.96 g × 3 × 188 g/mol / 296 g/mol = 5.64 g


Answer: 5.64 g of silver bromide